Taking the example of palmitic acid oxidation, explain the energetics of the fatty acid (5)
metabolism.
1
Expert's answer
2016-08-20T10:44:04-0400
The total oxidation of a fatty acid comprehends different processes:
1. – The activation of the fatty acid
2. – Beta-Oxidation
3. – Krebs Cycle.
For being metabolized, a fatty acid should experiment activation:
Fatty acid + CoA + ATP —-à Acyl CoA + AMP + 2(P)
The activation of the fatty acid requires 1 molecule of ATP, but since two energy rich bonds are hydrolyzed (the ATP is hydrolyzed to AMP and 2 (P) ) for energetic balance purposes it is considered that 2 ATP have been consumed in this activation process)

Palmitic acid has 16 carbons, so you can break it down into eight acetyl‐CoA molecules, with the formation of one FADH 2 and one NADH at each of the seven β‐oxidation steps. Electron transport starting with FADH 2 yields two ATPs and with NADH yields three ATPs. Therefore, β‐oxidation yields the equivalent of 35 ATPs per molecule of palmitic acid. Each acetyl‐CoA goes into the TCA cycle, where its metabolism yields three NADHs, one FAD, and one GTP directly, for a total of 12 ATPs. Thus, the ATPs produced are:
7 × 5 + 8 × 12 = 35 + 96 = 131
Two ATP equivalents were used to activate the fatty acid, leading to a total energy yield of 129 ATPs, over three times the amount of energy obtained from metabolizing a single molecule of glucose.